OpenStudy (anonymous):

What is the minimum number of bits required to represent 1786

OpenStudy (anonymous):

@jazydior : it would be bettr if askd in computer Science anyway let me try

OpenStudy (anonymous):

Another way of looking at the question is: What is the least number of digits needed to represent 1786 in base 2?

OpenStudy (anonymous):

2^0 = 1 2^1 = 2 2^2 = 4 ... ... keep going until you get a number bigger than 1786

OpenStudy (anonymous):

2^11 = 2048 So, that means you can write 1786 in base 2 using 11 digits

OpenStudy (anonymous):

1786 in base 2 (or in bits) is: 110 1111 1010 from http://www.cleavebooks.co.uk/scol/calnumba.htm

OpenStudy (anonymous):

Thank you all!